# Lowering PyTorch’s Memory Consumption for Selective Differentiation

### v1

@@ Abstract
Memory is a limiting resource for many deep learning tasks. Beside the neural network weights, one main memory consumer is the computation graph built up by automatic differentiation (AD) for backpropagation. We observe that PyTorch’s current AD implementation neglects information about parameter differentiability when storing the computation graph. This information is useful though to reduce memory whenever gradients are requested for a parameter subset, as is the case in many modern fine-tuning tasks. Specifically, inputs to layers that act linearly in their parameters (dense, convolution, or normalization layers) can be discarded whenever the parameters are marked as non-differentiable. We provide a drop-in, differentiability-agnostic implementation of such layers and demonstrate its ability to reduce memory without affecting run time [^1] Code:github.com/plutonium-239/memsave_torch..
@@ 1 Introduction & Motivation
The success of many deep learning applications is driven by scaling computational resources [32]. One important resource is GPU memory, specifically on low- and mid-end GPUs which usually offer between 6 to 16 GiB. Therefore, down-scaling the computational demands of deep learning is an important objective to widen its accessibility to researchers and practitioners with fewer hardware resources.
Two major memory consumers are the neural network weights, and the computation graph stored by the automatic differentiation (AD) engine. There exist various approaches to reduce them; e.g., the parameters can be compressed with low-precision data types (quantization [16, 19, 21]) or sparsified [12, 7], and the computation graph can be off-loaded to CPU [26], compressed [3], randomized [22], or partially recorded and re-computed (gradient checkpointing [10, 5]).
Figure 1: PyTorch’s AD is not agnostic to parameter differentiability. We consider a deep CNN that consists entirely of size-preserving convolutions and measure the forward pass’s peak memory when processing a mini-batch of size (256, 8, 256, 256), requiring 512 MiB storage. Memory increases linearly in the number of layers when all parameters are marked differentiable and remains constant when all parameters are marked non-differentiable. Surprisingly, when only one layer’s parameters are marked as differentiable the memory increases as if all subsequent parameters were marked differentiable. Our drop-in solution stores layer inputs depending on the layer parameter’s differentiability and reduces memory compared to the current PyTorch implementation.
Here, we consider a special AD scenario we call selective differentiation where gradients are requested only for a subset of the computation graph leafs (the neural network inputs and parameters). This approach has gained a lot of practical relevance in the era of large foundation models. For instance, fine-tuning techniques for pre-trained models rely on training only a subset of layers [36, 18], or extend some with parameter-efficient adapters [e.g. LoRA 15] which are then trained, keeping the original weights fixed. But it is also common in other applications such as generating adversarial examples [9] and neural style transfer [8] which optimize the input to a network with frozen weights.
We find that PyTorch’s [23] current AD allows for additional, simple, optimizations to further reduce memory consumption in the context of selective differentiation:
@@
    We observe that PyTorch’s AD neglects the differentiability of layer parameters when storing the computation graph (Figure 1). This information is useful though as it allows discarding inputs to linear layers whose parameters are marked as non-differentiable.
@@
    We provide a drop-in implementation of various layers that is agnostic to the parameter differentiability and demonstrate on various convolutional neural networks (CNNs) that it lowers on the default implementation’s memory footprint without increasing run time.
This easy-to-use insight benefits many tasks with selective differentiation and enables them to scale further. We hope it will stimulate further research into AD optimizations.
@@ 2 Selective Differentiation in PyTorch
PyTorch users can specify whether they want to compute gradients w.r.t. a tensor through its requires_grad attribute, which is dominantly inherited by child tensors: If any input (parent) to an operation is marked differentiable, the output (child) will also be differentiable and require the computation graph to be stored for backpropagation. As we will see, PyTorch does not check the differentiability of all parents, but rather starts storing the computation graph as if all parents were differentiable once it encounters a differentiable input. In the following example, we demonstrate this behavior, and how it misses out on possible optimizations for linear operations with non-differentiable parameters.
@@
We use PyTorch 2.2.1 and measure the forward pass’s peak memory of different neural networks as a proxy for the computation graph size stored by the AD engine. We chose peak memory as it is a reasonable proxy for the computation graph size and the relevant metric for causing out-of-memory errors in practise. On CPU, we use the memory_profiler library [24]. On GPU, we rely on torch.cuda.max_memory_allocated(), the maximum memory allocated by CUDA. Each measurement is performed in a separate Python session to avoid memory (de)allocation leaks between consecutive runs.
@@ 2.1 A Simple Example
Figure 2: PyTorch’s current behaviour of storing the computation graph. We illustrate this on a linear layer $\mathbf{z}=\mathbf{W}\mathbf{x}$. PyTorch stores the layer input whenever it is differentiable, although this is not necessary if the weight does not require gradients. Our approach uses this information to discard the layer input if possible.
Next, we observe the memory consumption in the context of selective differentiation. We consider two scenarios: In the first, all parameters after and including the fourth layer are differentiable, hence all layer inputs after the third layer must be stored in memory. In the second, only the fourth layer is differentiable and therefore only that layer’s input is necessary to compute gradients. However, we observe that both scenarios exhibit the same memory footprint (Figure 1, layers 4+ differentiable & layer 4 differentiable)!
We conclude that PyTorch stores a layer input with requires_grad = True although the layer’s parameters might be non-differentiable, and therefore not require it to be stored. This information can be useful to reduce the information stored for backpropagation, as we show with our implementation (Figure 1, layer 4 differentiable (ours)).
@@
Case All Input Conv Norm
(GPU) Default ResNet-101 7.82 (1.00) 7.82 (1.00) 7.82 (1.00) 7.82 (1.00)
+ swap Convolution 7.82 (1.00) 7.78 (0.99) 7.82 (1.00) 7.78 (0.99)
+ swap BatchNorm, ReLU 8.75 (1.12) 1.61 (0.21) 4.84 (0.62) 5.44 (0.70)
(CPU) Default ResNet-101 8.24 (1.00) 8.24 (1.00) 8.24 (1.00) 8.26 (1.00)
+ swap Convolution 8.35 (1.01) 8.30 (1.01) 8.33 (1.01) 8.26 (1.00)
+ swap BatchNorm, ReLU 9.37 (1.14) 2.52 (0.31) 5.53 (0.67) 6.37 (0.77)
@@
(GPU) Default ResNet-101 1.55 (1.00) 1.39 (0.90) 1.57 (1.01) 1.40 (0.90)
+ swap Convolution 1.50 (0.97) 1.37 (0.88) 1.50 (0.97) 1.36 (0.88)
+ swap BatchNorm, ReLU 1.46 (0.94) 1.32 (0.85) 1.45 (0.94) 1.30 (0.84)
(CPU) Default ResNet-101 22.78 (1.00) 16.81 (0.78) 22.84 (1.00) 16.66 (0.73)
+ swap Convolution 23.05 (1.01) 17.08 (0.75) 23.14 (1.02) 16.90 (0.74)
+ swap BatchNorm, ReLU 25.34 (1.11) 16.79 (0.74) 23.68 (1.04) 17.57 (0.77)
Table 1: Peak memory and run time comparison between PyTorch and our memory-saving layers on ResNet-101. Normalized values are relative to ‘All’ with PyTorch’s default layers.
@@
Table 2: Details of one residual block of ResNet101. Shapes of layer inputs (excluding the batch sizes) are shown on the right. The output of a ReLU layer is stored by both the activation layer itself, and the convolution layer it feeds into.
@@ 2.2 Implementation Details
2.2.1 Linear Layers
Consider a dense layer $\mathbf{z}=\mathbf{W}\mathbf{x}+\mathbf{b}$ with vector input $\mathbf{x}$ and output $\mathbf{z}$, weight matrix $\mathbf{W}$, and bias vector $\mathbf{b}$. Its input Jacobian depends on $\mathbf{W}$, the weight Jacobian on $\mathbf{x}$, and the bias Jacobian has no dependency [e.g. 6, Chapter 2.3]. During the forward pass, we check the differentiability of $\mathbf{W},\mathbf{x}$, and only store tensors required by the Jacobians that will be applied during backprogation. The same dependency pattern holds for other layers that process their inputs linearly w.r.t. their weight and add a bias term, such as convolution [see 2] and normalization [1, 33, 34, 17] layers. We implement them in exactly the same fashion.
@@ 2.2.2 Interaction with Other Layers
Real neural nets contain additional layers that are interleaved with linear layers: activations, dropout, pooling, etc. The information stored by such layers may overlap with the input to a linear layer, e.g. if an activation layer feeds into a convolution. Depending on the implementation details, our described optimizations might not apply in such cases because the linear layer’s input tensor could still be stored by the preceding activation. We encountered this effect for ReLU layers in real-world CNNs (Section 3), but were able to overcome it using a customized ReLU implementation that saves a boolean mask of the output instead. As PyTorch currently only supports 1-byte booleans, this leads to a 4x reduction of the stored tensor’s size, which could be further reduced to 32x with a 1-bit boolean implementation (which is in the works).
@@ 3 Real-World Examples
Case All Input Conv Norm All Input Conv Norm
DeepLabv3 (RN101) [4] 22.78 (1.00) 22.78 (1.00) 22.78 (1.00) 22.78 (1.00) 2.00 (1.00) 1.76 (0.88) 2.00 (1.00) 1.76 (0.88)
+ MemSave 24.90 (1.09) 4.28 (0.17) 13.72 (0.55) 15.17 (0.61) 1.97 (0.99) 1.76 (0.88) 1.99 (1.00) 1.77 (0.89)
EfficientNetv2-L [30, 31] 26.83 (1.00) 26.83 (1.00) 26.83 (1.00) 26.83 (1.00) 1.79 (1.00) 1.59 (0.89) 1.79 (1.00) 1.57 (0.88)
+ MemSave 26.83 (1.00) 10.45 (0.39) 18.62 (0.69) 18.64 (0.69) 1.66 (0.93) 1.45 (0.87) 1.63 (0.91) 1.45 (0.87)
FCN (RN101) [20] 22.19 (1.00) 22.19 (1.00) 22.19 (1.00) 22.19 (1.00) 1.94 (1.00) 1.72 (0.89) 1.94 (1.00) 1.71 (0.88)
+ MemSave 24.41 (1.10) 4.26 (0.17) 13.42 (0.55) 15.14 (0.62) 1.90 (0.98) 1.70 (0.88) 1.91 (0.98) 1.68 (0.87)
MobileNetv3-L [28, 14] 2.80 (1.00) 2.80 (1.00) 2.80 (1.00) 2.80 (1.00) 1.46 (1.00) 1.35 (0.92) 1.48 (1.01) 1.35 (0.92)
+ MemSave 2.94 (1.05) 0.87 (0.30) 1.90 (0.65) 1.91 (0.65) 1.44 (0.99) 1.30 (0.90) 1.44 (0.99) 1.31 (0.90)
ResNeXt101-64x4d [35] 15.18 (1.00) 15.18 (1.00) 15.18 (1.00) 15.18 (1.00) 1.71 (1.00) 1.53 (0.89) 1.68 (0.98) 1.53 (0.89)
+ MemSave 16.75 (1.10) 2.46 (0.15) 9.28 (0.55) 9.89 (0.59) 1.59 (0.93) 1.44 (0.84) 1.61 (0.94) 1.43 (0.84)
VGG-16 [29] 4.93 (1.00) 4.93 (1.00) 4.93 (1.00) N/A 1.50 (1.00) 1.33 (0.89) 1.49 (0.99) N/A
+ MemSave 4.30 (0.87) 2.60 (0.60) 4.30 (0.87) N/A 1.45 (0.97) 1.35 (0.90) 1.45 (0.97) N/A
Faster-RCNN (RN101) [27] 6.62 (1.00) 6.56 (0.99) 6.57 (0.99) 6.53 (0.99) 1.58 (1.00) 1.47 (0.93) 1.61 (1.02) 1.46 (0.92)
+ MemSave 7.21 (1.09) 1.90 (0.26) 4.55 (0.63) 4.43 (0.61) 1.51 (0.96) 1.39 (0.88) 1.54 (0.97) 1.38 (0.87)
SSDLite (MobileNetv3-L) [28] 0.53 (1.00) 0.53 (1.00) 0.53 (1.00) 0.53 (1.00) 1.54 (1.00) 1.39 (0.90) 1.53 (0.99) 1.40 (0.91)
+ MemSave 0.55 (1.04) 0.26 (0.48) 0.40 (0.73) 0.41 (0.74) 1.48 (0.96) 1.35 (0.88) 1.50 (0.97) 1.34 (0.88)
Table 3: Peak memory and run time comparison between PyTorch and our memory-saving layers on GPU on different scenarios and CNNs.
@@
Only the input to the neural net is differentiable. This situation resembles constructing adversarial examples [9] or style-transfers [8] by optimizing noisy inputs.
Conv:
Only convolution layers are differentiable. This situation is similar to surgical fine-tuning [18], which splits a network into different blocks, each containing a subset of layers, which are then trained one at a time.
@@
Only normalization layers are differentiable, resembling layer norm fine-tuning in LLMs [36].
We report GPU results on an NVIDIA RTX A6000 with 48 GiB of VRAM, and CPU results for an Apple M2 with 16 GiB of RAM. All networks are fed inputs of size (64, 3, 224, 224) and we measure according to the procedure described in Section 2. For object detection models, the batch size is 4 and 2 boxes are predicted per input image.

### v2

@@ Abstract
Memory is a limiting resource for many deep learning tasks. Beside the neural network weights, one main memory consumer is the computation graph built up by automatic differentiation (AD) for backpropagation. We observe that PyTorch’s current AD implementation sometimes neglects information about parameter differentiability when storing the computation graph. This information is useful though to reduce memory whenever gradients are requested for a parameter subset, as is the case in many modern fine-tuning tasks. Specifically, inputs to layers that act linearly in their parameters and inputs (fully-connected, convolution, or batch normalization layers in evaluation mode) can be discarded whenever the parameters are marked as non-differentiable. We provide a drop-in, differentiability-agnostic implementation of such layers111 Code and experiments available at github.com/plutonium-239/memsave_torch.and demonstrate its ability to reduce memory without affecting run time on popular convolution- and attention-based architectures.
@@ 1 Introduction & Motivation
The success of many deep learning applications is driven by scaling computational resources Thompson et al. (2020). One important resource is GPU memory, specifically on low- and mid-end GPUs which usually offer between 6 to 16 GiB. Therefore, down-scaling the computational demands of deep learning is an important objective to widen its accessibility to researchers and practitioners with fewer hardware resources. Two major memory consumers are the network weights, and the computation graph stored by the automatic differentiation (AD) engine. There exist various approaches to reduce their memory burden; e.g., the parameters can be compressed with low-precision data types (quantization Hubara et al. (2018); Li et al. (2017); Nagel et al. (2021)) or sparsified Hassibi & Stork (1992); Frantar & Alistarh (2022), and the computation graph can be off-loaded to CPU Ren et al. (2021), compressed Chen et al. (2021), randomized Oktay et al. (2021), or partially recorded and re-computed (gradient checkpointing Griewank & Walther (2008); Chen et al. (2016)).
Figure 1: PyTorch’s AD is sometimes not agnostic to parameter differentiability. We consider a deep CNN made of size-preserving convolutions and measure the forward pass’s peak memory when processing a mini-batch of size (256, 8, 256, 256), requiring 512 MiB memory. Memory increases linearly in the number of layers when all parameters are marked differentiable and remains constant when all parameters are marked non-differentiable. Surprisingly, when only one layer’s parameters are marked as differentiable the memory increases as if all subsequent parameters were marked differentiable. Our drop-in solution stores layer inputs depending on parameter differentiability and reduces memory compared to the current PyTorch implementation.
Here, we consider a special AD scenario we call selective differentiation where gradients are requested only for a subset of the computation graph leafs (the neural network inputs and parameters). This approach has gained a lot of practical relevance in the era of large foundation models. For instance, fine-tuning techniques for pre-trained models rely on training only a subset of layers Zhao et al. (2024); Lee et al. (2023); Zhu et al. (2024); Pan et al. (2024), or extend them with parameter-efficient adapters Hu et al. (2022) which are then trained, keeping the original weights fixed. But selective differentiation is also common in other applications such as generating adversarial examples Goodfellow et al. (2015) and neural style transfer Gatys et al. (2016) which optimize the input to a network with frozen weights.
We find that PyTorch’s Paszke et al. (2019) AD allows for additional, simple, optimizations to further reduce memory consumption in the context of selective differentiation:
@@
    We observe that PyTorch’s AD sometimes neglects the differentiability of layer parameters when storing the computation graph (Figure 1). This information is useful though as it allows discarding inputs to linear layers whose parameters are marked as non-differentiable.
@@
    We provide a drop-in implementation of various layers that is agnostic to the parameter differentiability and demonstrate on various convolutional neural networks (CNNs) and attention-based large language models (LLMs) that it lowers the default implementation’s memory footprint without increasing run time.
This easy-to-use insight benefits many tasks with selective differentiation and enables them to scale further. We hope it will stimulate future research into AD optimizations.
@@ 2 Selective Differentiation in PyTorch
PyTorch users can specify whether they want to compute gradients w.r.t. a tensor through its requires_grad attribute, which is dominantly inherited by child tensors: If any input (parent) to an operation is marked differentiable, the output (child) will also be differentiable and require the computation graph to be stored for backpropagation. As we will see, PyTorch does not always check the differentiability of all parents, but rather stores the computation graph as if all parents were differentiable once it encounters a differentiable input. In the following example, we demonstrate this behavior, and how it misses out on possible optimizations for linear operations with non-differentiable parameters.
@@
We use PyTorch 2.2.1 and measure the forward pass’s peak memory of different neural networks as a proxy for the computation graph size stored by the AD engine. We choose peak memory as it is a reasonable proxy for the computation graph size and the relevant metric for causing out-of-memory errors in practise. On GPU, we measure peak memory using torch.cuda.max_memory_allocated(), the maximum memory allocated by CUDA. To assess run time differences of our implementation, we compare a forward and backward pass with PyTorch. Each measurement is performed in a separate Python session to avoid memory (de)allocation leaks between consecutive runs.
@@ 2.1 A Simple Example
Figure 2: PyTorch’s behaviour of storing the computation graph, illustrated on a convolution ${\bm{Z}}={\bm{W}}*{\bm{X}}$. PyTorch stores the layer input whenever it is differentiable, although this is not necessary if the weight does not require gradients. Our approach uses this information to discard the layer input if possible. See Appendix C for computation graphs visualized with torchviz Zagoruyko et al. (2019).
Next, we observe the memory consumption in the context of selective differentiation. We consider two scenarios: In the first, all parameters after and including the fourth layer are differentiable, hence all layer inputs after the third layer must be stored in memory. In the second, only the fourth layer is differentiable and therefore only that layer’s input is necessary to compute gradients. However, we observe that both scenarios exhibit the same memory footprint (Figure 1, layers 4+ differentiable and layer 4 differentiable)!
We conclude that PyTorch stores a convolution layer’s input with requires_grad = True although the layer’s parameters might be non-differentiable. This information can be useful to reduce the information stored for backpropagation, as we show with our implementation (Figure 1, layer 4 differentiable (ours)).
In Appendix B, we repeat similar experiments with fully-connected, transpose convolution, and batch normalization layers. We find that PyTorch’s (transpose) convolutions and batch normalization layers in evaluation mode are not agnostic to the differentiability of their weights, irrespective of whether the model is compiled with torch.compile. Interestingly, fully-connected layers are already optimized.
@@
Case All Input Norm Surgical
Default ResNet-101 7.83 (1.00) 7.83 (1.00) 7.83 (1.00) 7.83 (1.00)
+ swap Convolution 7.83 (1.00) 7.78 (0.99) 7.78 (0.99) 7.83 (1.00)
+ swap ReLU 8.56 (1.09) 5.24 (0.67) 5.24 (0.67) 6.86 (0.88)
@@
Case All Input Norm Surgical
Default ResNet-101 0.45 (1.00) 0.37 (0.82) 0.37 (0.82) 0.41 (0.89)
+ swap Convolution 0.45 (0.98) 0.37 (0.80) 0.36 (0.80) 0.40 (0.87)
+ swap ReLU 0.44 (0.97) 0.36 (0.79) 0.36 (0.78) 0.41 (0.90)
Table 1: GPU peak memory and run time comparison between PyTorch and our memory-saving layers for ResNet-101. Normalized values are relative to ‘All’ with PyTorch’s default layers. We first swap only the convolution layers for our memory saving alternatives, which does not save memory (see main text for explanation). Then, we swap out other layers like ReLU, which significantly improves memory.
@@
Table 2: ResNet101’s residual block with layer input shapes, excluding batch size. The output of a ReLU is stored by both the activation layer and the convolution layer it feeds into.
@@ 2.2 Implementation Details
2.2.1 Convolution Layers
Consider a convolution layer ${\bm{Z}}={\bm{W}}*{\bm{X}}+{\bm{b}}$ with input ${\bm{X}}$, output ${\bm{Z}}$, kernel ${\bm{W}}$, and bias ${\bm{b}}$. Its input Jacobian depends on ${\bm{W}}$, the weight Jacobian on ${\bm{X}}$, and the bias Jacobian has no dependency (e.g. Dangel, 2023, Chapter 2.3). During the forward pass, we check the differentiability of ${\bm{W}},{\bm{X}}$, and only store tensors required by the Jacobians that will be applied during backpropagation. The same dependency pattern holds for other layers that process their inputs linearly w.r.t. their weight and input, and add a bias term, such as (transpose) convolution Chellapilla et al. (2006) and batch normalization Ioffe & Szegedy (2015) layers in evaluation mode. We implement them in exactly the same fashion.
@@ 2.2.2 Interaction with Other Layers
Real neural nets contain additional layers that are interleaved with linear layers: activations, dropout, pooling, etc. The information stored by such layers may overlap with the input to a linear layer, e.g. if an activation layer feeds into a convolution. Depending on the implementation details, our described optimizations might not apply in such cases because the linear layer’s input tensor could still be stored by the preceding activation. We encountered this effect for ReLU layers in real-world CNNs (Section 3), but were able to overcome it using a customized ReLU implementation that saves a boolean mask of the output. As PyTorch currently only supports 1-byte booleans, this leads to a 4x reduction of the stored tensor’s size, which could be further reduced to 32x (1-bit booleans are in the works). For dropout layers, we accept adding a slight computational overhead to avoid storing the output by saving the random state and re-generating the mask during backpropagation.
@@ 3 Real-World Examples
Case All Input Norm Surgical All Input Norm Surgical
DeepLabv3 (RN101) Chen et al. (2017) 22.82 (1.00) 22.82 (1.00) 22.82 (1.00) 22.82 (1.00) 0.93 (1.00) 0.73 (0.79) 0.73 (0.79) 0.76 (0.82)
+ MemSave 24.90 (1.09) 15.17 (0.66) 15.17 (0.66) 16.83 (0.74) 0.94 (1.01) 0.76 (0.82) 0.76 (0.81) 0.79 (0.85)
EfficientNetv2-L Tan & Le (2019; 2021) 26.81 (1.00) 26.81 (1.00) 26.81 (1.00) 26.81 (1.00) 0.77 (1.00) 0.62 (0.81) 0.62 (0.81) 0.68 (0.88)
+ MemSave 26.81 (1.00) 18.64 (0.70) 18.64 (0.70) 22.05 (0.82) 0.78 (1.02) 0.63 (0.82) 0.63 (0.82) 0.69 (0.90)
FCN (RN101) Long et al. (2015) 22.23 (1.00) 22.23 (1.00) 22.23 (1.00) 22.23 (1.00) 0.83 (1.00) 0.67 (0.80) 0.67 (0.80) 0.70 (0.84)
+ MemSave 24.39 (1.10) 15.15 (0.68) 15.15 (0.68) 16.80 (0.76) 0.87 (1.04) 0.70 (0.84) 0.69 (0.83) 0.74 (0.88)
Faster-RCNN (RN101) Ren et al. (2015) 6.84 (1.00) 6.84 (1.00) 6.84 (1.00) 6.84 (1.00) 0.77 (1.00) 0.66 (0.86) 0.66 (0.85) 0.69 (0.89)
+ MemSave 7.31 (1.07) 4.79 (0.70) 4.79 (0.70) 5.73 (0.84) 0.77 (0.99) 0.65 (0.84) 0.67 (0.86) 0.68 (0.88)
MobileNetv3-L Howard et al. (2019) 2.82 (1.00) 2.82 (1.00) 2.82 (1.00) 2.82 (1.00) 0.39 (1.00) 0.32 (0.82) 0.32 (0.82) 0.35 (0.89)
+ MemSave 2.96 (1.05) 1.91 (0.68) 1.91 (0.68) 2.52 (0.89) 0.40 (1.01) 0.32 (0.82) 0.32 (0.82) 0.35 (0.89)
ResNeXt101-64x4d Xie et al. (2017) 15.15 (1.00) 15.15 (1.00) 15.15 (1.00) 15.15 (1.00) 0.65 (1.00) 0.53 (0.82) 0.53 (0.82) 0.58 (0.90)
+ MemSave 16.75 (1.11) 9.87 (0.65) 9.87 (0.65) 13.32 (0.88) 0.64 (0.98) 0.52 (0.80) 0.52 (0.80) 0.56 (0.87)
SSDLite (MobileNetv3-L) Sandler et al. (2018) 0.54 (1.00) 0.53 (0.97) 0.53 (0.97) 0.53 (0.97) 0.63 (1.00) 0.59 (0.92) 0.55 (0.87) 0.58 (0.91)
+ MemSave 0.57 (1.04) 0.41 (0.75) 0.41 (0.75) 0.50 (0.92) 0.66 (1.05) 0.57 (0.90) 0.55 (0.86) 0.58 (0.92)
VGG-16 Simonyan & Zisserman (2015) 4.93 (1.00) 4.93 (1.00) N/A 5.05 (1.02) 0.37 (1.00) 0.28 (0.77) N/A 0.31 (0.84)
+ MemSave 4.30 (0.87) 3.08 (0.62) N/A 3.15 (0.63) 0.38 (1.05) 0.30 (0.82) N/A 0.33 (0.89)
Table 3: GPU peak memory and run time comparison between PyTorch and our memory-saving layers for CNNs.
@@
Only the input to the neural net is differentiable. This situation resembles constructing adversarial examples Goodfellow et al. (2015) or style-transfers Gatys et al. (2016) by optimizing noisy inputs.
Surgical:
Only the first quarter of layers are differentiable. This situation is similar to surgical fine-tuning Lee et al. (2023), which splits a network into different blocks, each containing a subset of layers, to be trained one at a time.
@@
Only normalization layers are differentiable, resembling layer norm fine-tuning in LLMs Zhao et al. (2024).
We report GPU results on an NVIDIA RTX A6000 with 48 GiB of VRAM. All CNNs are fed inputs of size (64, 3, 224, 224) and we measure according to the procedure described in Section 2. For object detection models, the batch size is 4 and 2 boxes are predicted per input image.